Three Rods A,B.C of the same length and same cross sectional area are joined their thermal conductivities are in ratio 1:2:1.5 if open end of Aand C at 200degree celcius and 18 degree celcius the temperature at the fuction of Aand B in equilibrium

Grade:11

2 Answers

karan
4 years ago
3/k(s) = 1/k + 1/2k + 1/3k

∴ k(s) = 18/13 k

( Θ / t ) combined = ( Θ / t ) A

18/13 k x A (200 - 18)/ 3l = k A (200 - Θ) / l

∴ Θ = 116°

Vikas TU
4 years ago
DEAR STUDENT 
Kindly take the last ratio in k as 2/3 , Because in the question , it is written that it is 1:2:1.5 .
Next part you can continue with the solution , 
Good Luck